OpenStudy (anonymous):

Simplify. √108 - 3√8 -√27

OpenStudy (anonymous):

3√3 - 5√2 -3√3 - 5√2 3√3 - 6√2 -3√3 - 6√2

OpenStudy (anonymous):

3 sqrt(3)-6 sqrt(2)

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Trust me it is the answer

OpenStudy (anonymous):

i understand that but i want to know how to get it

OpenStudy (anonymous):

? simplify

OpenStudy (anonymous):

you mind showing me the work?

OpenStudy (anonymous):

To tired to just know thats the answer

OpenStudy (anonymous):

then why are you one this site? The sites code of conduct states that you are supposed to help the person through the answer not just give it to them

OpenStudy (anonymous):

on*

OpenStudy (anonymous):

any person will tell you thats the answer

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

You first simplify each radical, if possible. \[\sqrt{108} = 6\sqrt{3}\]\[-3\sqrt{8} = -6\sqrt{2}\]\[-1\sqrt{27} = -3\sqrt{3}\]

OpenStudy (anonymous):

So you problem now becomes \[6\sqrt{3} - 6\sqrt{2} - 3\sqrt{3}\] Remember you can only combine like radicals, so \[6\sqrt{3} - 3\sqrt{3} =?\]

OpenStudy (anonymous):

(The difference of 6sqrt(3) and 3sqrt(3)) - 6sqrt(2) If you need any further explaination feel free to ask